### v2

@@ Abstract
A graph neural network (GNN) approach is introduced in this work which enables mesh-based three-dimensional super-resolution of fluid flows. In this framework, the GNN is designed to operate not on the full mesh-based field at once, but on localized meshes of elements (or cells) directly. To facilitate mesh-based GNN representations in a manner similar to spectral (or finite) element discretizations, a baseline GNN layer (termed a message passing layer, which updates local node properties) is modified to account for synchronization of coincident graph nodes, rendering compatibility with commonly used element-based mesh connectivities. The architecture is multiscale in nature, and is comprised of a combination of coarse-scale and fine-scale message passing layer sequences (termed processors) separated by a graph unpooling layer. The coarse-scale processor embeds a query element (alongside a set number of neighboring coarse elements) into a single latent graph representation using coarse-scale synchronized message passing over the element neighborhood, and the fine-scale processor leverages additional message passing operations on this latent graph to correct for interpolation errors. Demonstration studies are performed using hexahedral mesh-based data from Taylor-Green Vortex flow simulations at Reynolds numbers of 1600 and 3200. Through analysis of both global and local errors, the results ultimately show how the GNN is able to produce accurate super-resolved fields compared to targets in both coarse-scale and multiscale model configurations.
@@
Graph neural networks , Super-resolution , Fluid dynamics , Taylor-Green Vortex , Deep learning
@@ 1 Introduction
Development of GNN-based models in the fluid dynamics community has largely centered around surrogate forecasting [40, 46, 43, 47], sub-grid modeling [8, 7], and unstructured autoencoding applications [44, 48]. Little focus has been placed on leveraging the capabilities of GNN-based architectures for super-resolution of fluid flows, especially in three spatial dimensions. An exception can be found in Ref. [49], which applies GNNs for reconstruction in two spatial dimensions; however, this work focuses on reconstruction of flow-fields from a set of sensor measurements, and does not explicitly move from coarse to fine grid representations. The goal of the present work is to fill this gap and develop a novel multiscale GNN architecture for the mesh-based super-resolution task applied to arbitrary three-dimensional (3D) mesh-based discretizations. The modeling framework, inspired by the recent physics-informed super-resolution work of Ref. [50], operates within a localized patch-based approach and is conceptually in-line with p-refinement strategies leveraged in finite element simulations [51]. The specific contributions of this work are as follows:
@@
    A novel GNN architecture is implemented for mesh-based 3D super-resolution of fluid flows, which takes as input a graph representation of unstructured flow data stored on a spectral/finite element-discretized CFD mesh.
@@
    To facilitate mesh-based GNN representations in a manner similar to spectral element discretizations, the baseline message passing layer, popularized for fluid flows in Ref. [40], is modified here to account for synchronization of coincident graph nodes, rendering compatibility with commonly used finite element-based mesh connectivities.
@@
    Demonstration studies are performed using a Taylor-Green Vortex 3D configuration at Reynolds numbers of 1600 and 3200, with analysis conducted using time- and Reynolds-extrapolated data. The domain is discretized using unstructured hexahedral mesh data structures consistent with those utilized in complex geometry finite element simulations. In other words, although TGV mesh used here is regular with equally-sized elements, the data are arranged in an unstructured format (there is no structured grid/connectivity assumption being made). As such, the methods introduced here can be readily applied to complex-geometry meshes.
@@
2 Description of Dataset
Training and evaluation datasets in this work are generated by high-fidelity CFD simulations of the Taylor-Green Vortex (TGV). The TGV is a classic benchmark problem in the CFD literature, with early numerical studies performed by Orszag [52] and Brachet et al [53]. It has since been used in a wide variety of applications related to the numerical simulation of inviscid and turbulent fluid flows; recent use-cases include, but are not limited to: (a) spurring the development of low-dissipation numerical schemes for flow solvers in complex geometries [54], (b) usage as a benchmark configuration to evaluate CFD solver numerics [55, 56, 57], and (c) facilitating the development of super-resolution and turbulence models [13, 58, 59], as done here.
The governing equations of TGV flow are the incompressible Navier-Stokes (NS) equations, given by
@@ 2 Description of Dataset
In the above equations, ${\bf u}={\bf u}({\bf x},t)$ represents the three-dimensional fluid velocity defined at a particular point in physical space ${\bf x}$ and time $t$, the quantity $p=p({\bf x},t)$ is the scalar pressure, and Re denotes the Reynolds number, a critical parameter that sets both the turbulence intensity and degree of scale separation observed in the fluid flow. To observe complex turbulent flow with appreciable scale separation in the TGV configuration, one must set $\text{Re}=\mathcal{O}(1000)$ or higher [53]. As such, in this work, two values for Reynolds number are considered to generate training and evaluation datasets: $\text{Re}=1600$ and $\text{Re}=3200$.
$u_{x}({\bf x},0)$ $=\text{sin}(x)\text{cos}(y)\text{cos}(z),$ (3)
$u_{y}({\bf x},0)$ $=-\text{cos}(x)\text{sin}(y)\text{cos}(z),$ (4)
$u_{z}({\bf x},0)$ $=0,$ (5)
@@
with $x,y,z\in[-\pi,\pi]$, satisfying periodic boundary conditions. The initial condition is effectively two-dimensional – it is characterized completely by large length scales through a single low-frequency Fourier mode, as visualized by the vorticity contours in Fig. 1(a). Numerical simulations of the TGV configuration are performed here with NekRS [60], an extensively verified open-source GPU-based exascale CFD code developed at Argonne National Laboratory. NekRS solves the governing equations using a spectral element discretization in space on unstructured meshes, which can be composed of elements consisting of wedges, tetrahedra, and hexahedra (a regular mesh composed of hexahedral elements is used in this work) [61] Interested readers are directed to Ref. [60] for further details of solver numerics implementation, time-stepping schemes, and scalability.
@@
At a particular Reynolds number, the flow solver generates a high-dimensional simulation trajectory describing the time-evolution of the TGV velocity field on a set of spatial discretization points. These discretization points are prescribed by a mesh. Similar to conventional mesh-based simulation strategies (e.g., as in finite volume and finite element discretizations), in the spectral element approach of NekRS, the mesh is represented by a series of non-overlapping elements. A schematic of the mesh used for the TGV simulations is shown in Fig. 1(b). In the solution procedure, the velocity field is solved on a set of quadrature points within (and at the interface of) each element. Examples of individual elements extracted from the TGV mesh are shown in Fig. 1(c) – the figure illustrates how, for the same underlying mesh, different levels of fidelity can be achieved through prescription of a different number of quadrature points within each element. NekRS leverages a Guass-Lobatto-Legendre (GLL) quadrature [60], resulting in non-uniformly arranged spatial points within each element for high-order discretizations. The fidelity of the quadrature is characterized by the polynomial order, $P$, resulting in $N_{p}=(P+1)^{3}$ points per element in 3D. Shown in Fig. 1(c) are two such polynomial order settings: (1) $P=1$, which corresponds to the coarse flow-fields resulting in 8 GLL points per element, all of which are coincident with the mesh vertices, and (2) $P=7$, which represents fine flow-fields resulting in 512 GLL points per element.
@@
$\mathbf{Y}(t)=\left[\,\begin{array}[]{c}{\bf y}_{1}(t)\\ \hline\cr{\bf y}_{2}(t)\\ \hline\cr\vdots\\ \hline\cr{\bf y}_{N_{e}}(t)\\ \end{array}\,\right]\in\mathbb{R}^{(N_{e}N_{p})\times 3},\quad{\bf y}_{i}=\left[\begin{array}[]{c}{\bf u}_{i}({\bf x}_{1},t)\\ \vdots\\ {\bf u}_{i}({\bf x}_{N_{p}},t)\\ \end{array}\right]\in\mathbb{R}^{N_{p}\times 3},\quad i=1,\ldots,N_{e}.$ (6)
@@
In Eq. 6, ${\bf Y}(t)$ is the full flow-field at time $t$, consisting of a total of $N_{e}N_{p}$ discretization points, where $N_{e}$ is the total number of mesh elements and $N_{p}$ is the total number of quadrature points per element. As shown in Fig. 1(b), the total number of mesh elements is fixed to $36^{3}$, resulting in roughly 24M discretization points on the fine $P=7$ mesh [^1] An element polynomial order of $P=7$ on this mesh is roughly equivalent to a conventional $256^{3}$ structured grid discretization, which is deemed to satisfy the requirements for fully resolved TGV simulations at the two Reynolds numbers considered here [53]. and 0.37M points on the coarse $P=1$ mesh. Per Eq. 6, each flow-field ${\bf Y}(t)$ is composed of a concatenation of element-local flow-fields, where ${\bf y}_{i}$ denotes the velocity field in the $i$-th element. Distinctions between flow-field representations at different polynomial orders on the same mesh are indicated by subscripts where appropriate: a $P=7$ flow-field is denoted ${\bf Y}_{7}(t)$ (with the element-local quantity given by ${\bf y}_{7,i}(t)$). It should be noted that the flow-field representation in Eq. 6, while consistent with the spectral element discretization employed in NekRS, also can be readily generalized to other mesh-based discretizations such as those found in finite-volume and discontinuous Galerkin methods (the finite-volume case, for example, results in $N_{p}=1$).
@@
Considering the above discussion, the datasets used in the training and evaluation procedure are sourced from temporal ranges characterized by turbulent flow behavior near the peak dissipation rate for both Reynolds numbers. To facilitate a supervised training strategy (to be described in detail in Sec. 3), a coarse-fine flow-field pairing at a particular time instant $t$ is generated through a coarsening operation on a set of target P=7 flow-fields, producing a coarsened counterpart on the P=1 mesh (referred to as the projected DNS solution). This coarsening operation is expressed by a linear restriction operator ${\cal R}$, such that ${\bf Y}_{1}(t)={\cal R}{\bf Y}_{7}(t)$. When moving from $P>1$ to $P=1$, the operator $\cal R$ in the spectral element framework coincides with an indexing operation on the flow-fields, since the $P=1$ element GLL points are a subset of the $P>1$ GLL points. However, without loss of generality, $\cal R$ can be interpreted as a general non-invertible filter acting on an instantaneous flow-field [10].
@@
Visualization of a fine-to-coarse flow-field pair is shown in Fig. 3(a) and (b) for P=7 and P=1, respectively, for a $\text{Re}=1600$ snapshot. Consistent with expectations, the coarsening operation retains many of the large-scale flow characteristics (such as flow symmetry about the x-y quadrants); however, the elimination of fine-scale flow-features is evident, posing a challenging super-resolution task. This elimination is highlighted in the corresponding energy spectrum curves in Fig. 3(c). The spectra explicitly highlight the elimination of high-wavenumber energy content through the coarsening procedure for both Reynolds numbers, as well as increase in high-energy content in the target flow-fields with Reynolds number. Ultimately, training data consists of three snapshots at $t=[8,9,10]$, resulting in a total of $139968$ coarse-fine element pairs in the training set for each Reynolds number.

### v3

@@ Abstract
A graph neural network (GNN) approach is introduced in this work which enables mesh-based three-dimensional super-resolution of fluid flows. In this framework, the GNN is designed to operate not on the full mesh-based field at once, but on localized meshes of elements (or cells) directly. To facilitate mesh-based GNN representations in a manner similar to spectral (or finite) element discretizations, a baseline GNN layer (termed a message passing layer, which updates local node properties) is modified to account for synchronization of coincident graph nodes, rendering compatibility with commonly used element-based mesh connectivities. The architecture is multiscale in nature, and is comprised of a combination of coarse-scale and fine-scale message passing layer sequences (termed processors) separated by a graph unpooling layer. The coarse-scale processor embeds a query element (alongside a set number of neighboring coarse elements) into a single latent graph representation using coarse-scale synchronized message passing over the element neighborhood, and the fine-scale processor leverages additional message passing operations on this latent graph to correct for interpolation errors. Demonstration studies are performed using hexahedral mesh-based data from Taylor-Green Vortex and backward-facing step flow simulations at Reynolds numbers of 1600 and 3200. Through analysis of both global and local errors, the results ultimately show how the GNN is able to produce accurate super-resolved fields compared to targets in both coarse-scale and multiscale model configurations. Reconstruction errors for fixed architectures were found to increase in proportion to the Reynolds number. Geometry extrapolation studies on a separate cavity flow configuration show promising cross-mesh capabilities of the super-resolution strategy.
@@
Graph neural networks , Super-resolution , Fluid dynamics , Taylor-Green Vortex , Backward-facing step , Deep learning
@@ 1 Introduction
Development of GNN-based models in the fluid dynamics community has largely centered around surrogate forecasting [40, 46, 43, 47], sub-grid modeling [8, 7], and unstructured autoencoding applications [44, 48]. Little focus has been placed on leveraging the capabilities of GNN-based architectures for super-resolution of fluid flows, especially in three spatial dimensions. An exception can be found in Ref. [49], which applies GNNs for reconstruction in two spatial dimensions; however, this work focuses on reconstruction of flow-fields from a set of sensor measurements, and does not explicitly move from coarse to fine grid representations. The goal of the present work is to fill this gap and develop a novel multiscale GNN architecture for the mesh-based super-resolution task applied to arbitrary three-dimensional (3D) mesh-based discretizations. The modeling framework, inspired by the recent physics-informed super-resolution work of Ref. [50], operates within a localized patch-based approach and is conceptually in-line with p-refinement strategies leveraged in finite element simulations [51]. The contributions of this work are as follows:
@@
    A new multiscale graph neural network architecture is introduced to handle mesh-based super-resolution. Specifically, coarse-scale and fine-scale message passing operations, separated by an intermediary unpooling layer, are used to recover the fine-scale flow-field given a coarse-scale flow-field as input.
@@
    Building on previous work [50, 52], an element-local model form is adopted, which enables the construction of a training set using a relatively small number of snapshots. The element-local formulation used in this work is mesh-based, leveraging graph representations of mesh neighborhoods of spectral elements.
@@
    To facilitate mesh-based graph representations in a manner similar to spectral element discretizations, the baseline message passing layer, popularized for fluid flows in Ref. [40], is modified here to account for synchronization of coincident graph nodes, rendering compatibility with commonly used finite element-based mesh connectivities.
  • 5.
    Demonstrations are performed on complex mesh-based physical configurations in three spatial dimensions. Specifically, super-resolution capability is independently assessed for a Taylor-Green Vortex and Backward-Facing Step flow at Reynolds numbers of 1600 and 3200, and a separate Cavity flow configuration is used to demonstrate geometry extrapolation capability of trained models.
@@
2 Description of Datasets
Training and evaluation datasets in this work are generated from high-fidelity computational fluid dynamics (CFD) simulations in three different configurations: (1) the Taylor-Green Vortex (TGV), (2) flow over a backward-facing step (BFS), and (3) flow over a cavity. Before detailing the datasets for each configuration, the governing equations, simulation procedure, and solution representation are first described.
2.1 Simulation Procedure
For all three configurations, the governing equations are the incompressible Navier-Stokes (NS) equations in three spatial dimensions (with isothermal and Newtonian assumptions and no body force), given by
@@ 2.1 Simulation Procedure
In the above equations, ${\bf u}={\bf u}({\bf x},t)$ represents a three-dimensional fluid velocity defined at a particular point in physical space ${\bf x}$ and time $t$, the quantity $p=p({\bf x},t)$ is the scalar pressure, and Re denotes the Reynolds number, a critical parameter that sets both the turbulence intensity and degree of scale separation observed in the fluid flow. To observe complex turbulent flow with appreciable scale separation in the above mentioned configurations, one must set Re to appreciably high values. As such, in this work, two values for Reynolds number are considered to generate training and evaluation datasets: $\text{Re}=1600$ and $\text{Re}=3200$. In the case of the BFS and cavity, these Reynolds numbers are defined with respect to the step and cavity heights, respectively.
@@
Numerical simulations are performed here with NekRS [53], an extensively verified open-source GPU-based exascale CFD code developed at Argonne National Laboratory. NekRS solves the governing equations using a spectral element discretization in space on unstructured meshes, which can be composed of elements consisting of wedges, tetrahedra, and hexahedra (a regular mesh composed of hexahedral elements is used in this work) [54]. Interested readers are directed to Ref. [53] for further details of solver numerics implementation, time-stepping schemes, and scalability.
@@
At a particular Reynolds number, the flow solver generates a high-dimensional simulation trajectory describing the time-evolution of a velocity field on a set of spatial discretization points. These discretization points are prescribed by a mesh. Similar to conventional mesh-based simulation strategies (e.g., as in finite volume and finite element discretizations), in the spectral element method formulation of NekRS, the mesh is represented by a series of non-overlapping elements. In the solution procedure, the velocity field is solved on a set of quadrature points within (and at the interface of) each element. Examples of individual elements extracted from the TGV mesh (the same also applies to BFS and cavity meshes) are shown in Fig. 1(c) – the figure illustrates how, for the same underlying mesh, different levels of fidelity can be achieved through prescription of a different number of quadrature points within each element. NekRS leverages a Guass-Lobatto-Legendre (GLL) quadrature [53], resulting in non-uniformly arranged spatial points within each element for high-order discretizations. The fidelity of the quadrature is characterized by the polynomial order, $P$, resulting in $N_{p}=(P+1)^{3}$ points per element in 3D. Shown in Fig. 1(c) are two such polynomial order settings: (1) $P=1$, which corresponds to the coarse flow-fields resulting in 8 GLL points per element, all of which are coincident with the mesh vertices, and (2) $P=7$, which represents fine flow-fields resulting in 512 GLL points per element.
@@
$\mathbf{Y}(t)=\left[\,\begin{array}[]{c}{\bf y}_{1}(t)\\ \hline\cr{\bf y}_{2}(t)\\ \hline\cr\vdots\\ \hline\cr{\bf y}_{N_{e}}(t)\\ \end{array}\,\right]\in\mathbb{R}^{(N_{e}N_{p})\times 3},\quad{\bf y}_{i}(t)=\left[\begin{array}[]{c}{\bf u}_{i}({\bf x}_{1},t)\\ \vdots\\ {\bf u}_{i}({\bf x}_{N_{p}},t)\\ \end{array}\right]\in\mathbb{R}^{N_{p}\times 3},\quad i=1,\ldots,N_{e}.$ (3)
@@
In Eq. 3, ${\bf Y}(t)$ is the full flow-field at time $t$, consisting of a total of $N_{e}N_{p}$ discretization points, where $N_{e}$ is the total number of mesh elements and $N_{p}$ is the total number of quadrature points per element. Per Eq. 3, each flow-field ${\bf Y}(t)$ is composed of a concatenation of element-local flow-fields, where ${\bf y}_{i}(t)$ denotes the velocity field in the $i$-th element. Distinctions between flow-field representations at different polynomial orders on the same mesh are indicated by subscripts where appropriate: a P=7 flow-field is denoted ${\bf Y}_{7}(t)$ (with the element-local quantity given by ${\bf y}_{7,i}(t)$). It should be noted that the flow-field representation in Eq. 3, while consistent with the spectral element discretization employed in NekRS, also can be readily generalized to other mesh-based discretizations such as those found in finite-volume and discontinuous Galerkin methods (the finite-volume case, for example, results in $N_{p}=1$).
@@
In all configurations described below, to facilitate a supervised training strategy (to be described in detail in Sec. 3), a coarse-fine flow-field pairing at a particular time instant $t$ is generated through a coarsening operation on a set of target P=7 flow-fields, producing a coarsened counterpart on the P=1 mesh (referred to as the projected target solution). This coarsening operation is expressed by a linear restriction operator ${\cal R}$, such that ${\bf Y}_{1}(t)={\cal R}{\bf Y}_{7}(t)$. When moving from $P>1$ to $P=1$, the operator $\cal R$ in the spectral element framework coincides with an indexing operation on the flow-fields, since the $P=1$ element GLL points are a subset of the $P>1$ GLL points. However, without loss of generality, $\cal R$ can be interpreted as a general non-invertible filter acting on an instantaneous flow-field [10].
2.2 Taylor-Green Vortex (TGV)
The TGV is a classic benchmark problem in the CFD literature, with early numerical studies performed by Orszag [55] and Brachet et al [56]. It has since been used in a wide variety of applications related to the numerical simulation of inviscid and turbulent fluid flows; recent use-cases include, but are not limited to: (a) spurring the development of low-dissipation numerical schemes for flow solvers in complex geometries [57], (b) usage as a benchmark configuration to evaluate CFD solver numerics [58, 59, 60], and (c) facilitating the development of super-resolution and turbulence models [13, 61, 62], as done here.
The TGV initial condition is shown in Fig. 1(a), and a schematic of the mesh used for the TGV simulation is shown in Fig. 1(b). The total number of mesh elements is fixed to $36^{3}$, resulting in roughly 24M discretization points on the fine P=7 mesh [^1] An element polynomial order of $P=7$ on this mesh is roughly equivalent to a conventional $256^{3}$ structured grid discretization, which is deemed to satisfy the requirements for fully resolved TGV simulations at the two Reynolds numbers considered here [56]. and 0.37M points on the coarse P=1 mesh. A description of the physical characteristics of the TGV is provided in A for the unfamiliar reader. Ultimately, the datasets used in the training and evaluation procedure are sourced from temporal ranges characterized by turbulent flow behavior near the peak dissipation rate for both Re=1600 and 3200.
@@ 2.2 Taylor-Green Vortex (TGV)
Visualization of a fine-to-coarse flow-field pair is shown in Fig. 2(a) and (b) for P=7 and P=1, respectively, for a $\text{Re}=1600$ snapshot. Consistent with expectations, the coarsening operation retains many of the large-scale flow characteristics (such as flow symmetry about the x-y quadrants); however, the elimination of fine-scale flow-features is evident, posing a challenging super-resolution task. This elimination is highlighted in the corresponding energy spectrum curves in Fig. 2(c). The spectra explicitly highlight the elimination of high-wavenumber energy content through the coarsening procedure for both Reynolds numbers, as well as increase in high-energy content in the target flow-fields with Reynolds number. Training data for the TGV consists of three snapshots at $t=[8,9,10]$ time units, resulting in a total of $139968$ coarse-fine element pairs in the training set for each Reynolds number.
@@
2.3 Backward-Facing Step (BFS) and Cavity
The BFS and cavity mesh configurations are shown in Fig. 3. These geometries have been studied in a number of experimental [63, 64, 65], numerical [66, 67], and machine learning [68, 69] works in the context of physical analysis and modeling of geometry-induced separated turbulent flows. In both cases, flow enters from the inflow (detailed in white arrows in Fig. 3) into a channel upon encountering a backward-facing step. In both the BFS and cavity, the step anchor triggers flow separation, resulting in a shear layer that then reattaches to the wall further downstream (in the case of the cavity, reattachment may also occur on the opposing vertical wall). The separation triggers a fully turbulent flow, with complex multiscale flow features downstream of the separation point, until the flow exits at the domain outlet. At higher Reynolds numbers, alongside increased turbulent intensity, flow reattachment points and other large-scale dynamical quantities of interest (such as recirculation and shear layer dynamics) become noticeably more complex.
The boundary conditions for both cases include fixed inflow and stabilized outflow [70] settings, span-wise periodicity, and no-slip walls everywhere else, with the exception of an initial ramp-up region that extends along the top and bottom walls for a streamwise distance of $5h$ from the inlet (shown in the 2D schematics in Fig. 3). In this ramp-up region, slip wall boundary conditions are used (instead of no-slip) to mitigate discontinuity effects associated with a fixed-velocity inflow. Ultimately, in both BFS and cavity cases, the flow traverses a streamwise distance of $10h$ before encountering the step anchor – the first half of this distance is the ramp-up phase, with no-slip conditions prescribed afterwards.
Figure 3: Meshes for backward-facing step (left) and cavity (right) configurations. Bottom schematics show 2D cross-sections.
The BFS and cavity meshes consist of a total of 46480 and 42440 hexahedral elements-per-snapshot, respectively. It should be noted that the P=7 target cases here are under-resolved relative to direct numerical simulation requirements (the resolution here is roughly equivalent to $h/28$, where $h$ is the step/cavity height). Despite this, for this study, the target resolution is adequate for capturing highly complex turbulent flow features in both configurations.
This is confirmed through visualization of fine-coarse flow-field pairs, as is shown in Fig. 4. Comparison of Fig. 4(a) and (b) highlights the added complexity of increasing Reynolds number from 1600 to 3200 through the addition of smaller length-scales near and downstream of the reattachment point. Comparison of Fig. 4(a) and (c) highlights the physical differences between BFS and cavity flow-fields at the same Reynolds number; the primary difference is due to flow impingement on the rear wall of the cavity (a reverse-step), in which additional recirculation zones are present in the place of a clean reattachment point to the bottom wall.
Figure 4: (a) Coarse-fine training snapshot pairs for BFS configuration at Re=1600, showing z-component vorticity contours (top), x-y plane slices of vorticity (z-component) taken at z=1 (middle), and x-y plane slices of velocity magnitude (bottom). (b) Same as (a), but for BFS at Re=3200. (c) Same as (a), but for cavity at Re=1600. Vorticity contours show $w_{z}=-5$ in red and $w_{z}=5$ in blue. Domains are cropped in x- and y-directions for near-step visualization.
Training data for BFS and cavity configurations are sourced from a set of snapshots after $t=100$ time units of simulation, after which all of the initial vortex transients are shed. After this point, for each of the BFS and cavity cases, 10 training snapshots are extracted, each separated by $\Delta t=2.5$ time units. From each of these snapshots, elements to populate the training dataset are sampled according to the following setup (illustrated in the lower schematics in Fig. 3): (a) 100% of elements below $y=h$ (i.e., below one step-height above the step anchor) are retained, (b) 100% of elements above $y=4.5$ are retained, and (c) 10% of the remaining elements (largely in the freestream) are randomly sampled. This implicitly biases the training algorithm away from the freestream, and towards the boundary layer and turbulent flow features.
After this procedure, the total number of training elements is 138060 for the BFS configuration and 103320 for the cavity, resulting in similar total training element counts for all three configurations. Lastly, it is emphasized that the scope of the BFS and cavity configurations are constrained to the following: (a) the BFS is leveraged as an additional demonstration case for GNN super-resolution performance, in that additional GNNs are trained and evaluated using data from this configuration; (b) the cavity is leveraged as a geometry extrapolation testbed: both TGV- and BFS-trained models are evaluated on the cavity to determine geometry generalization capability of models trained from physically different data sources.